# All-Electron Molecular Tunnel Ionization Based on the Weak-Field Asymptotic Theory in the Integral Representation

### v1

@@ IV.3 OCS and CO
We first look at OCS, whose HOMO has a $\pi$ symmetry and is doubly degenerate. In our simulations, we choose the O–C and C–S bond lengths to be $1.123$ and $1.564$, respectively. The double-peak structure of $|G_{00}|^{2}$ obtained by OE-WFAT seen in Fig. 2(a) is a manifestation of the $\pi$ symmetry of the HOMO. The structure factors obtained by ME-WFAT (Fig. 2(b)) still retain the same general features as OE-WFAT (Fig. 2(a)) for both parabolic channels simulated here but this time the magnitude of the peak of the $G_{00}$ channel around $\beta\sim 125^{\circ}$ is three times larger than the other peak. Here, we will show that this difference is largely due to the different treatment of the dipole moment in the two methods. We run a OE-WFAT simulation but this time, the origin and dipole moment are chosen according to ME-WFAT prescription. This means that the origin is such that $\mathbf{D}^{(N-1)}=0$ and the dipole moment is given in the explanation immediately following Eq. (38). The rest, such as the orbital and integral formulas, are the same as those used to obtain Fig. 2(a). The structure factors resulting from these modifications to OE-WFAT are shown in Fig. 2(c). We see that in the case of OCS, mixing the OE-WFAT prescription for orbital and integral formula with the ME-WFAT prescription for the origin and dipole moment closely recovers the shape of the structure factors obtained by ME-WFAT. We note that the angle-dependent ionization yield for OCS obtained using DFT-based IR ME-WFAT has been calculated in Ref. [30], and it agrees with $|G_{00}|^{2}$ in Fig. 2(b).
@@
Next, we turn to CO whose HOMO has a $\sigma$ symmetry. We choose the C–O bond length to be $1.102$. The OE-WFAT and ME-WFAT structure factors are shown in Fig. 3(a) and 3(b), respectively, and we see a contrasting behavior in the location of the global maximum. IR OE-WFAT predicts that the ionization should be maximum at $\beta=180^{\circ}$ (as does the TR OE-WFAT in Ref. [13, 10]), that is, when the field points from O to C, while ME-WFAT predicts the opposite. It is interesting to note that a recent calculation on CO has also been done addressing the effect of the induced multielectron polarization (MEP) [42] on the ionization yield. One of the findings in [42] is that without MEP, the angle-resolved ion yield from CO looks very similar to $|G_{00}|^{2}$ in Fig. 3(a), whereas with MEP, the ion yield resembles closely $|G_{00}|^{2}$ of Fig. 3(b). Experiments [43] as well as time-dependent Hartree-Fock [44] calculations also produce angle-resolved ionization of CO in which the ionization is most favorable when the field points from C to O, and their shape is very similar to the yield with MEP in Ref. [42] and $|G_{00}|^{2}$ in Fig. 3(b). What is interesting here is that the IR ME-WFAT introduced in this work does not capture MEP effects (which may be done by including the 1st order correction of ME-WFAT).
@@ IV.4 N2, HCOOH, and CH3F
In the previous section, we demonstrated that for some molecules, ME-WFAT structure factors differ from the OE-WFAT counterpart mostly by the value of the dipole moment used in the exponential factor of Eq. (III.3). In these cases, ME-WFAT structure factors may be recovered by deliberately using the ME-WFAT dipole moment and origin in the OE-WFAT formulation. In this section, we will show that ME-WFAT and OE-WFAT calculations can differ much less trivially when the Dyson orbital connecting the neutral and cation ground states notably differ from the neutral HOMO.
@@
It is well-known that the occupied molecular orbitals of N2 calculated by HF have an incorrect ordering—experiments suggested that the first ionization from N2 should produce a $\Sigma$ cation state, while Koopman’s theorem says that the first ionization should leave the cation in a $\Pi$ state [45, 46, 41]. See Ref. [41] for an experimental orientation-dependent ionization yield of N2. In the context of OE-WFAT where the ionizing orbital must be chosen manually, the HF simulation for N2 can lead to the wrong angle-dependent ionization yield because the $\pi$ character of the false HOMO will obviously dominate since it has the lowest binding energy. Using a N–N bond length of $1.066$, we run OE-WFAT calculation N2 and obtain Fig. 4(a) where we see that OE-WFAT $|G_{00}|^{2}$ calculated using the HOMO of N2 exhibits a $\pi_{u}$ symmetry. Fig. 4(b) shows the results of ME-WFAT, where $|G_{00}|^{2}$ is seen to have the correct $\sigma$ symmetry. The structure factors shown in Fig. 4(b) also look very similar to those obtained by OE-WFAT using HOMO-1 of N2 [10]. The reason of the success of ME-WFAT in this case is that the Dyson orbital, being a linear combination of the occupied molecular orbitals of the neutral, has the largest component (largest $|\mathcal{P}(i)|$) that corresponds to HOMO-1 (see Table 3). In a sense, the overlap integral between the neutral and cation states defining the Dyson orbital automatically selects HOMO-1 as the orbital that is closest to the actual ionizing orbital. This result also illustrates how a properly relaxed cation wave function can facilitate the construction of an accurate ionizing orbital.
@@
Figure 5: The structure factors $|G_{00}|^{2}$ of HCOOH calculated using (a) OE-WFAT and (b) ME-WFAT. (c) The shape of HCOOH. (d) The HOMO used to calculate the $|G_{00}|^{2}$ in (a). (e) The Dyson orbital used to calculate the $|G_{00}|^{2}$ in (b). (f) The values of $\mathcal{P}^{\prime}(i)=(-1)^{i}\mathcal{P}(i)$ of several molecular orbitals with the largest contribution to the Dyson orbital. The sign symbol above each bar indicates its sign. The orbital picture shown above the sign symbol is the shape of the corresponding neutral orbital.
@@
Figure 6: (a) and (b): $|G_{00}|^{2}$ of CH3F obtained by OE-WFAT using HOMO(–) and HOMO(+), respectively. (c) and (d): $|G_{00}|^{2}$ of CH3F obtained by ME-WFAT using Dyson(–) and Dyson(+), respectively. (e) The shape of CH3F. (f) HOMO(–). (g) Dyson(–). (h) The values of $\mathcal{P}^{\prime}(i)$ of several molecular orbitals with the largest contribution to Dyson(–). The above structure factors are based on the reference that $(\beta,\gamma)=(0^{\circ},0^{\circ})$ happens when the C–F bond coincides the $z$-axis.
@@
Next, we examine HCOOH (formic acid), a planar molecule. The atomic coordinates in angstrom when $(\beta,\gamma)=(0^{\circ},0^{\circ})$ are C: $(0,0,0.415)$, O: $(0,-1.011,-0.437)$, O: $(0,1.139,0.110)$, H: $(0,-0.372,1.444)$, and H: $(0,-0.658,-1.322)$. According to Ref. [47], formic acid is a good candidate to study the difference between using the Dyson orbital and the HOMO to calculate the ionization rate due to the notable contribution of occupied orbitals other than the HOMO in the linear combination of Eq. (59). $|G_{00}|^{2}$ resulting from an ionization from the HOMO calculated with OE-WFAT is shown in Fig. 5(a). It features a single lobe centered around $(\beta,\gamma)=(67^{\circ},90^{\circ})$. The ME-WFAT version of this structure factor is shown in Fig. 5(b), where we see that the single lobe seen in Fig. 5(a) persists. The most notable difference with OE-WFAT is the appearance of a second, weaker lobe centered around $(\beta,\gamma)=(128^{\circ},270^{\circ})$ and the ’leakage’ of the main lobe toward $\beta=0^{\circ}$. These features, which are absent in the OE-WFAT structure factor, indicate that the there are some occupied orbitals of the neutral that interfere with the dominating HOMO in forming the ionizing orbital. Fig. 5(d) and 5(e) compare the shapes of HOMO and Dyson orbital, respectively, with Fig. 5(f) presenting the individual contribution of some of the most important occupied orbitals of the neutral in terms of their coefficient, $\mathcal{P}(i)$, in the linear combination of Eq. (59). We see that the contribution of HOMO-5 is not so small.
@@
As the last example, we examine CH3F, a molecule having an equilibrium geometry that possesses a $C_{3v}$ symmetry, where the axis of the three-fold rotational symmetry coincides with the CF bond. The C–F and C–H bond lengths are $1.360$ and $1.080$, respectively, with the HCF angle of $108.93^{\circ}$. Its HOMO is doubly degenerate and can be identified based on their reflection symmetry with respect to the plane containing one of the three FCH planes, here, we denote the symmetric and antisymmetric functions as HOMO(+) and HOMO(–), respectively. We choose the reflection plane to be the $xz$-plane. Fig. 6(a) and 6(b) are the $|G_{00}|^{2}$ structure factors obtained by OE-WFAT using HOMO(–) and HOMO(+), respectively. The ground state of CH3F+ is also doubly degenerate, which is the consequence of the equal probabilities of removing an electron from either HOMO(+) or HOMO(–) of the neutral to form the ground state cation. This means that one may construct two ”degenerate” Dyson orbitals, Dyson(+) and Dyson(–). The ME-WFAT $|G_{00}|^{2}$ structure factors using Dyson(–) and Dyson(+) are shown in Fig. 6(c) and 6(d), respectively. We see that the positions as well as the relative intensities are visibly different compared to OE-WFAT results (Fig. 6(a) and (b)), for example, the peaks are at smaller $\beta$ for ME-WFAT structure factors (Fig. 6(c) and 6(d)) and the middle bottom lobe in Fig. 6(d) (ME-WFAT) is at a lower intensity than the corresponding lobe in Fig. 6(b) (OE-WFAT). Again, we can understand this observation by looking at the shapes of the HOMO and Dyson orbital shown in Fig. 6(f) and (g), respectively, for the odd symmetry. The balancing of the upper and lower pairs of lobes of HOMO(–) (Fig. 6(f)) in the formation of the Dyson orbital (Fig. 6(g)) are due to the appreciable interference between HOMO(–) and HOMO-2(–) (HOMO-2 is also doubly degenerate) that has a comparable magnitude of $\mathcal{P}(i)$ (see Fig. 6(h)).
@@ V Conclusions and Outlook
The main factors that can make the angular dependence of the ionization yield calculated by OE-WFAT and ME-WFAT noticeably differ are (i) the ionizing orbital (the Dyson orbital in ME-WFAT and a molecular orbital in OE-WFAT), (ii) the dipole moments, or a combination of both. The structure factors calculated by OE-WFAT and ME-WFAT are in very good agreement if the Dyson orbital is close to the HOMO and the HOMO dipole moment is similar to the difference between total initial and final state’s dipole moment. An example of such molecule is O2. When the difference in (ii) is more prominent, new features will appear in ME-WFAT structure factors not seen in OE-WFAT (N2, HCOOH, and CH3F are examples in this category), which include producing a totally different shape for the structure factor. From our analysis on N2 (Section IV.4), we see that ME-WFAT has the property such that it tends to select the most appropriate ionizing orbital having the right symmetry as verified by experiments.
One obvious direction for improvement in the formulation we have presented is to include the first order correction to $f_{\nu\sigma}^{n^{\prime}n}$ to simulate the effect of the distortion of the wave functions on the ionization probability – this has been done in the TR ME-WFAT [35]. The inclusion of this induced effect will modify the effective dipole moment of the molecule, which will in turn affect the net force that drives the ionized electron. Another prospect is the use of multi-configuration wave functions, which has been found to pose a problem for TR ME-WFAT [26]. In particular, IR ME-WFAT can be used to accurately calculate ionization from molecules dominated by multireference configurations, such as those containing transition metals, provided accurate multireference wave functions are provided. Apart from the possibility of studying correlation effects in tunnel ionization, ME-WFAT with multi-configuration wave functions also enables one to ionize from an excited initial state. This is expected to become increasingly important as field strength increases as more and more excited levels of the neutral will get populated.

### v2

@@ IV.3 OCS and CO
We first look at OCS, whose HOMO has a $\pi$ symmetry and is doubly degenerate. OCS is one of the most frequently studied molecules for its strong-field ionization properties [72, 73, 38, 53, 39]. In our simulations, we choose the O–C and C–S bond lengths to be $1.123$ and $1.564$, respectively. The double-peak structure of $|G_{00}|^{2}$ obtained by OE-WFAT seen in Fig. 2(a) is a manifestation of the $\pi$ symmetry of the HOMO. The structure factors obtained by ME-WFAT (Fig. 2(b)) still retain the same general features as OE-WFAT (Fig. 2(a)) for both parabolic channels simulated here but this time the magnitude of the peak of the $G_{00}$ channel around $\beta\sim 125^{\circ}$ is three times larger than the other peak. Here, we will show that this difference is largely due to the different treatment of the dipole moment in the two methods. We run a OE-WFAT simulation but this time, the origin and dipole moment are chosen according to ME-WFAT prescription. This means that the origin is such that $\mathbf{D}^{(N-1)}=0$ and the dipole moment is given in the explanation immediately following Eq. (38). The rest, such as the orbital and integral formulas, are the same as those used to obtain Fig. 2(a). The structure factors resulting from these modifications to OE-WFAT are shown in Fig. 2(c). We see that in the case of OCS, mixing the OE-WFAT prescription for orbital and integral formula with the ME-WFAT prescription for the origin and dipole moment closely recovers the shape of the structure factors obtained by ME-WFAT.
We note that alignment-dependent rates, rather than orientation-dependent ones, of OCS seem to be more commonly calculated or measured [74, 72, 73, 38]. In these alignment-dependent curves, the maximum occurs at $\beta=90^{\circ}$. Approximately emulating an alignment-dependent rate from our orientation-dependent rate $|G_{00}|^{2}$ in Fig. 2(b), however, produces a pair of peaks at $\beta=38^{\circ}$ and $\beta=142^{\circ}$ instead. The absence of a maximum at $\beta=90^{\circ}$ in our result is due to insufficient suppression of the secondary peak at $\beta=28^{\circ}$ in the orientation-dependent curve of Fig. 2(b) and the global maximum in the same figure occurring at a slightly too large of an angle $\beta$. Despite this discrepancy in the alignment-dependent curve, our orientation-dependent curves of OCS calculated through IR-MEWFAT (Fig. 2(b)) and DFT-based IR-MEWFAT [53] agree with RT-TDDFT result (see Ref. [53]). In our opinion, a comparison of orientation-dependent rates, rather than alignment-dependent ones, should offer a higher confidence level since weak structures in the latter tend to get washed out due to the incoherent sum of rates from opposite orientations.
@@
Next, we turn to CO whose HOMO has a $\sigma$ symmetry. We choose the C–O bond length to be $1.102$. The OE-WFAT and ME-WFAT structure factors are shown in Fig. 3(a) and 3(b), respectively, and we see a contrasting behavior in the location of the global maximum. IR OE-WFAT predicts that the ionization should be maximum at $\beta=180^{\circ}$ (as does the TR OE-WFAT in Ref. [17, 14]), that is, when the field points from O to C, while ME-WFAT predicts the opposite. It is interesting to note that a recent calculation on CO has also been done addressing the effect of the field-induced MEP [75] on ionization yields. One of the findings in Ref. [75] is that without MEP, the angle-resolved ion yield from CO looks very similar to $|G_{00}|^{2}$ in Fig. 3(a), whereas with MEP, the ion yield resembles closely $|G_{00}|^{2}$ of Fig. 3(b). Experiments [76] as well as time-dependent Hartree-Fock [77] calculations also produce angle-resolved ionization of CO in which the ionization is most favorable when the field points from C to O, and their shape is very similar to the yield with MEP in Ref. [75] and $|G_{00}|^{2}$ in Fig. 3(b). What is interesting here is that the IR ME-WFAT introduced in this work does not capture MEP effects.
@@ IV.4 N2, HCOOH, and CH3F
In the previous section, we demonstrated that for some molecules, ME-WFAT structure factors differ from the OE-WFAT counterpart mostly by the value of the dipole moment used in the exponential factor of Eq. (III.3). In these cases, ME-WFAT structure factors may be recovered by deliberately using the ME-WFAT dipole moment and origin in the OE-WFAT formulation. We will now show that ME-WFAT and OE-WFAT calculations can differ much less trivially when the Dyson orbital connecting the neutral and cation ground states notably differ from the neutral HOMO.
@@
It is well-known that the occupied MO of N2 calculated by HF have an incorrect ordering—experiments suggested that the first ionization from N2 should produce a $\Sigma$ cation state, while Koopman’s theorem says that the first ionization should leave the cation in a $\Pi$ state [78, 79, 70]. See Ref. [70] for an experimental orientation-dependent ionization yield of N2. In the context of OE-WFAT where the ionizing orbital must be chosen manually, the HF simulation for N2 can lead to the wrong angle-dependent ionization yield because the $\pi$ character of the false HOMO will obviously dominate since it has the lowest binding energy. Using a N–N bond length of $1.066$, we run OE-WFAT calculation N2 and obtain Fig. 4(a) where we see that OE-WFAT $|G_{00}|^{2}$ calculated using the HOMO of N2 exhibits a $\pi_{u}$ symmetry. Fig. 4(b) shows the results of ME-WFAT, where $|G_{00}|^{2}$ is seen to have the correct $\sigma$ symmetry. The structure factors shown in Fig. 4(b) also look very similar to those obtained by OE-WFAT using HOMO-1 of N2 [14]. The reason of the success of ME-WFAT in this case is that the Dyson orbital, being a linear combination of the occupied MO of the neutral, has the largest component (largest $|\mathcal{P}(i)|$) that corresponds to HOMO-1 (see Table 3). In a sense, the overlap integral between the neutral and cation states defining the Dyson orbital automatically selects HOMO-1 as the orbital that is closest to the actual ionizing orbital. This result also illustrates how a properly relaxed cation wave function can facilitate the construction of an accurate ionizing orbital.
@@
Figure 5: Structure factors $|G_{00}|^{2}$ from (a) HOMO, (c) HOMO-4, and (d) HOMO-5 of HCOOH calculated using OE-WFAT. The magnitude of $|G_{00}|^{2}$’s in panel (c) and (d) have been divided by $100$. (b) $|G_{00}|^{2}$ calculated by ME-WFAT. (e) Image of the HCOOH molecule as well as the $z-$axis which determines the $(0^{\circ},0^{\circ})$ direction. (f) The HOMO used to calculate the $|G_{00}|^{2}$ in (a). (g) The Dyson orbital used to calculate the $|G_{00}|^{2}$ in (b). (h) The values of $\mathcal{P}^{\prime}(i)=(-1)^{i}\mathcal{P}(i)$ of several MO with the largest contribution to the Dyson orbital. The sign symbol above each bar indicates its sign. The orbital picture shown above the sign symbol is the shape of the corresponding neutral orbital.
@@
Figure 6: (a) and (b): $|G_{00}|^{2}$ of CH3F obtained by OE-WFAT using HOMO(–) and HOMO(+), respectively. (c) and (d): $|G_{00}|^{2}$ of CH3F obtained by ME-WFAT using Dyson(–) and Dyson(+), respectively. (e) The shape of CH3F. (f) HOMO(–). (g) Dyson(–). (h) The values of $\mathcal{P}^{\prime}(i)$ of several MO with the largest contribution to Dyson(–). The above structure factors are based on the reference that $(\beta,\gamma)=(0^{\circ},0^{\circ})$ happens when the C–F bond coincides the $z$-axis.
@@
Next, we examine HCOOH (formic acid), a planar molecule (see Fig. 5(e)). The atomic coordinates in angstrom when $(\beta,\gamma)=(0^{\circ},0^{\circ})$ are C: $(0,0,0.415)$, O: $(0,-1.011,-0.437)$, O: $(0,1.139,0.110)$, H: $(0,-0.372,1.444)$, and H: $(0,-0.658,-1.322)$. According to Ref. [80], formic acid is a good candidate to study the difference between using the Dyson orbital and the HOMO to calculate the ionization rate due to the notable contribution of occupied orbitals other than the HOMO in the linear combination of Eq. (59). $|G_{00}|^{2}$ resulting from an ionization from the HOMO calculated with OE-WFAT is shown in Fig. 5(a). It features a single lobe centered around $(\beta,\gamma)=(67^{\circ},90^{\circ})$. The ME-WFAT version of this structure factor is shown in Fig. 5(b), where we see that the single lobe seen in Fig. 5(a) persists. The most notable difference with OE-WFAT is the appearance of a second, weaker lobe centered around $(\beta,\gamma)=(128^{\circ},270^{\circ})$ and the ’leakage’ of the main lobe toward $\beta=0^{\circ}$. These features, which are absent in the OE-WFAT structure factor, indicate that there are some occupied orbitals of the neutral that interfere with the dominating HOMO in forming the ionizing orbital. Fig. 5(f) and 5(g) compare the shapes of HOMO and Dyson orbital, respectively, with Fig. 5(h) presenting the individual contribution of some of the most important occupied orbitals of the neutral in terms of their coefficient, $\mathcal{P}(i)$, in the linear combination of Eq. (59).
We see that the contribution of HOMO-5 is not so small compared to the largest contribution coming from HOMO. The contribution from HOMO-5 to the Dyson orbital is most likely also responsible for the sharp difference in the magnitude of $|G_{00}|^{2}$ calculated by OE-WFAT (Fig. 5(a)) and ME-WFAT (Fig. 5(b)). This can be seen from OE-WFAT structure factor from HOMO-5 shown in Fig. 5(d), where we see that it has a maximum at about the same orientation as the structure factor in Fig. 5(a)  [^1] For this analysis, only the angular dependence of OE-WFAT-calculated structure factors from HOMO-4 and HOMO-5 are relevant. The magnitude, however, is irrelevant because it is determined mainly by orbital energies, which are not present in the formulation of ME-WFAT.. This may suggest that the regions of HOMO and HOMO-5 that are responsible for their respective orientation of maximum ionization rate interfere in a way that drastically reduces the maximum ionization probability from the Dyson orbital. There is another MO that produces maximum ionization at about the same angle as HOMO and HOMO-5, namely HOMO-4, whose OE-WFAT structure factor is shown in Fig. 5(c). However, its corresponding $\mathcal{P}(i)$ is significantly smaller than that of HOMO-5.
@@
As the last example, we examine CH3F, a molecule having an equilibrium geometry that possesses a $C_{3v}$ symmetry, where the axis of the three-fold rotational symmetry coincides with the CF bond. The C–F and C–H bond lengths are $1.360$ and $1.080$, respectively, with the HCF angle of $108.93^{\circ}$. Its HOMO is doubly degenerate and can be identified based on their reflection symmetry with respect to the plane containing one of the three FCH planes, here, we denote the symmetric and antisymmetric functions as HOMO(+) and HOMO(–), respectively. We choose the reflection plane to be the $xz$-plane. Fig. 6(a) and 6(b) are the $|G_{00}|^{2}$ structure factors obtained by OE-WFAT using HOMO(–) and HOMO(+), respectively. The ground state of CH3F+ is also doubly degenerate, which is the consequence of the equal probabilities of removing an electron from either HOMO(+) or HOMO(–) of the neutral to form the ground state cation. This means that one may construct two “degenerate” Dyson orbitals, Dyson(+) and Dyson(–). The ME-WFAT $|G_{00}|^{2}$ structure factors using Dyson(–) and Dyson(+) are shown in Fig. 6(c) and 6(d), respectively. We see that the positions as well as the relative intensities are visibly different compared to OE-WFAT results (Fig. 6(a) and (b)), for example, the peaks are at smaller $\beta$ for ME-WFAT structure factors (Fig. 6(c) and 6(d)) and the middle bottom lobe in Fig. 6(d) (ME-WFAT) is at a lower intensity than the corresponding lobe in Fig. 6(b) (OE-WFAT). Again, we can understand this observation by looking at the shapes of the HOMO and Dyson orbital shown in Fig. 6(f) and (g), respectively, for the odd symmetry. The balancing of the upper and lower pairs of lobes of HOMO(–) (Fig. 6(f)) in the formation of the Dyson orbital (Fig. 6(g)) are due to the appreciable interference between HOMO(–) and HOMO-2(–) (HOMO-2 is also doubly degenerate) that has a comparable magnitude of $\mathcal{P}(i)$ (see Fig. 6(h)).
@@
As a final note regarding the Dyson orbital, the $\mathcal{P}(i)$ coefficients, which encode the contribution of the neutral molecule’s MO to the Dyson orbital, is independent of the orbital energies of these MO. This is worth mentioning because one might be tempted to think that $|\mathcal{P}(i)|$ should die down monotonically as a function of orbital energy difference relative to that of the orbital that dominates the calculated ionization channel (usually the HOMO for ground-state-to-ground-state channels). This is an incorrect view because these contributions only tell us how much of these MO is needed to build an optimal ionizing orbital (the Dyson orbital). The correct way to compare ionization probabilities of a group of MO in terms of their orbital energies is by looking at the ionization channels that can be approximated by the removal of an electron from the said MO where their ionization potentials are approximately calculated as an orbital energy of this MO.
@@ V Conclusions and Outlook
The main factors that can make the angular dependence of the ionization yield calculated by OE-WFAT and ME-WFAT noticeably differ are (i) the ionizing orbital (the Dyson orbital in ME-WFAT and a MO in OE-WFAT), (ii) the dipole moments, or a combination of both. The structure factors calculated by OE-WFAT and ME-WFAT are in very good agreement if the Dyson orbital is close to the HOMO and the HOMO dipole moment is similar to the difference between total initial and final state’s dipole moment. An example of such molecule is O2. When the difference in (ii) is more prominent, new features will appear in ME-WFAT structure factors not seen in OE-WFAT (N2, HCOOH, and CH3F are examples in this category), which include producing a totally different shape for the structure factor. From our analysis on N2 (Section IV.4), we see that ME-WFAT has the property such that it tends to select the most appropriate ionizing orbital having the right symmetry as verified by experiments.
One obvious direction for improvement in the formulation we have presented is to include the first order correction to $f_{\nu\sigma}^{n^{\prime}n}$ to simulate the effect of the distortion of the wave functions on the ionization probability – this has been done in the TR ME-WFAT [62]. The inclusion of this induced effect will modify the effective dipole moment of the molecule, which will in turn affect the net force that drives the ionized electron. Another prospect is the use of multiconfiguration wave functions, which has been found to pose a problem for TR ME-WFAT [49]. In particular, IR ME-WFAT can be used to accurately calculate ionization from molecules dominated by multireference configurations, such as those containing transition metals, provided accurate multireference wave functions are provided. Apart from the possibility of studying correlation effects in tunnel ionization, ME-WFAT with multiconfiguration wave functions also enables one to ionize from an excited initial state. This is expected to become increasingly important as field strength increases as more and more excited levels of the neutral will get populated.